OpenStudy (anonymous):

Which is the simplified form of 15 times q times r to the fifth power all over 3 times q cubed times r to the fourth power.? 5 times r all over q squared. 12 times r all over q squared. 5 times q squared all over r. 12 times q squared all over r.

jimthompson5910 (jim_thompson5910):

so the problem is \[\large \frac{15qr^{5}}{3qr^{4}}\] right?

OpenStudy (anonymous):

thats very close but on the denomenator there is a power of 3 on the q but other whys it's perfect.

jimthompson5910 (jim_thompson5910):

oh my bad

jimthompson5910 (jim_thompson5910):

so \[\large \frac{15qr^{5}}{3q^{3}r^{4}}\] ??

OpenStudy (anonymous):

yes sir,

jimthompson5910 (jim_thompson5910):

ok thx, one sec

OpenStudy (anonymous):

My problem is i know how to simplify the 15 and 3 but i don't know how to simplify the exponents.

jimthompson5910 (jim_thompson5910):

Here's how you simplify \[\large \frac{15qr^{5}}{3q^{3}r^{4}}\] \[\large \frac{15}{3}*\frac{q}{q^{3}}*\frac{r^{5}}{r^{4}}\] \[\large \frac{5}{1}*\frac{q^{1-3}}{1}*\frac{r^{5-4}}{1}\] \[\large \frac{5}{1}*\frac{q^{-2}}{1}*\frac{r^1}{1}\] \[\large \frac{5}{1}*\frac{1}{q^{2}}*\frac{r}{1}\] \[\large \frac{5*1*r}{1*q^{2}*1}\] \[\large \frac{5r}{q^{2}}\]

jimthompson5910 (jim_thompson5910):

when dividing exponential terms, you subtract the exponents

jimthompson5910 (jim_thompson5910):

so for instance x^7 divided by x^3 = x^(7-3) = x^4

OpenStudy (anonymous):

ok so thats how Q^3 became q

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

well like you said you subtract the exponents when, your dividing so would you divide the bottom exponents but the alike exponents on the top. which is actullay my bad because the q^3 - q^1 = q^2?

jimthompson5910 (jim_thompson5910):

yes it's the other way around actually q divided by q^3 q^1 divided by q^3 q^(1-3) q^(-2) 1/( q^2 )

jimthompson5910 (jim_thompson5910):

I'm using the idea that x^( -k ) = 1/( x^k )

OpenStudy (anonymous):

what does the 1/ in front mean?

jimthompson5910 (jim_thompson5910):

1 over

jimthompson5910 (jim_thompson5910):

1/x means \(\large \frac{1}{x}\)

jimthompson5910 (jim_thompson5910):

so instead of saying "2 over 3", you would say 2/3

OpenStudy (anonymous):

oh wait is it like how you have a negative in a equation and you have to, take the negative out with a positive?

jimthompson5910 (jim_thompson5910):

kind of...

jimthompson5910 (jim_thompson5910):

the negative exponent is made positive by flipping the fraction

OpenStudy (anonymous):

oh ok, sorry if im being a little oblivious.

jimthompson5910 (jim_thompson5910):

no its ok, it's good you're asking

OpenStudy (anonymous):

well thankyou sir,

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Im gonna go practice this on some other questions, if i have any difficulties I will just post another question, but again thank you for your help.

jimthompson5910 (jim_thompson5910):

sure thing, and good idea, practice makes perfect